ANS 3. Hans car can travel 162 miles on 6 gallons of gas. How far can Hans travell if he uses 11 gallons of gas.
Answer #1

applying the unitary method to solve this -

162 miles on 6 gallons of gas = x miles on 11 gallons of gas

i.e. 162/6 = x/11

x = ( 162 * 11 ) / 6

x = 297 miles.

Han's car will travel 297 miles on 11 gallons of gas.